OpenStudy (anonymous):

need help with these two Factoring questions, Factoring fully: 1. (m+n)^2-m-n and 2. (a^2-13a)^2-900

jimthompson5910 (jim_thompson5910):

# 1 (m+n)^2-m-n (m+n)^2-1(m+n) Do you see what to do next?

OpenStudy (anonymous):

oh ya i see so i take the (m+n) as a common value and then put the rest in a seperate value! thanks so much jim! do u think u could help me out with the second one as well?

jimthompson5910 (jim_thompson5910):

exactly, you factor out the GCF m+n to go from (m+n)^2-1(m+n) to (m+n)(m+n-1) and you're done with #1

jimthompson5910 (jim_thompson5910):

# 2 (a^2-13a)^2-900 is the same as (a^2-13a)^2-(30)^2 since 30^2 = 900

jimthompson5910 (jim_thompson5910):

now use the difference of squares factoring rule

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

okay so i'm now using the 30^2 instead of the 900?

jimthompson5910 (jim_thompson5910):

yes and you're using this rule a^2 - b^2 = (a-b)(a+b)

jimthompson5910 (jim_thompson5910):

well i should use different variables x^2 - y^2 = (x-y)(x+y)

jimthompson5910 (jim_thompson5910):

in this case, x = a^2-13a and y = 30

jimthompson5910 (jim_thompson5910):

now tell me what you get when you factor using this rule

OpenStudy (anonymous):

okay just trying it out now

OpenStudy (anonymous):

i got part of the answer correct but the other part is not correct, the answer i got was (a-9) (a-4) (a-10) (a+3) but i should have gotten (a-10) (a-3) (a+2) (a-15) what am i doing wrong in this case?

jimthompson5910 (jim_thompson5910):

when you factor (a^2-13a)^2-(30)^2 using the difference of squares rule, you should get (a^2-13a-30)(a^2-13a+30)

jimthompson5910 (jim_thompson5910):

then you factor a^2-13a-30 to get (a-15)(a+2) notice how -15 plus 2 = -13 and -15 times 2 = -30

jimthompson5910 (jim_thompson5910):

and you factor a^2-13a+30 to get (a-10)(a-3) notice how -10 plus -3 = -13 and -10 times -3 = 30

jimthompson5910 (jim_thompson5910):

so that's why (a^2-13a-30)(a^2-13a+30) factors completely into (a-15)(a+2)(a-10)(a-3)

OpenStudy (anonymous):

ohh okay thanks a lot i got the answer now thank you for your help :)

jimthompson5910 (jim_thompson5910):

you're welcome